40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 23 21:30:29 2009, Jenne, Update, LSC, Measured MC length 
    Reply  Tue Nov 24 08:00:16 2009, rana, Update, LSC, Measured MC length 
       Reply  Tue Nov 24 10:36:21 2009, Koji, Update, LSC, Measured MC length 
          Reply  Wed Nov 25 03:05:15 2009, rob, Update, Locking, Measured MC length 
             Reply  Wed Nov 25 14:29:08 2009, rob, Update, Locking, Measured MC length--FSS trend FSStrendpowerjump.png
                Reply  Wed Nov 25 16:13:27 2009, rana, Update, PSL, Measured MC length--FSS trend 
                   Reply  Wed Nov 25 16:44:52 2009, Koji, Update, PSL, Measured MC length--FSS trend 091125_FSS.png
             Reply  Wed Nov 25 15:38:08 2009, rob, Update, Locking, Measured MC length 
Message ID: 2325     Entry time: Wed Nov 25 03:05:15 2009     In reply to: 2320     Reply to this: 2332   2333
Author: rob 
Type: Update 
Category: Locking 
Subject: Measured MC length 

Quote:

What I meant was the VCO driver, not the FSS box.

As for the frequency, all written numbers were the Marconi displays.
The number on the frequency counter was also recorded, and so will be added to the previous entry shortly... 

Quote:

I propose that from now on, we indicate in the elog what frequencies we're referring to. In this case, I guess its the front panel readback and not the frequency counter -- what is the frequency counter readback? And is everything still locked to the 10 MHz from the GPS locked Rubidium clock?

Plus, what FSS Box? The TTFSS servo box? Or the VCO driver? As far as I know, the RC trans PD doesn't go through the FSS boxes, and so its a real change. I guess that a bad contact in the FSS could have made a huge locking offset.

 

 

Locking has gone sour.  The CARM to MCL handoff, which is fairly early in the full procedure and usally robust, is failing reliably. 

As soon as the SUS-MC2_MCL gain is reduced, lock is broken.  There appears to be an instability around 10Hz.  Not sure if it's related.

ELOG V3.1.3-